OpenStudy (moonlight93):

can you simplify this.. x=cuberoot of (2y+3) x= 3√(2y+3)

OpenStudy (anonymous):

there isn't anything else you can do unless you want to solve for y

OpenStudy (anonymous):

simplify from x to y y=(1/18)(x^2-27)

OpenStudy (moonlight93):

i want to solve for y @sourwing

OpenStudy (moonlight93):

@BabiiMuffin can you show me how, please

OpenStudy (fibonaccichick666):

So, what do you have to do to "undo" a cube root?

OpenStudy (moonlight93):

raise by power of 3 on both sides? @FibonacciChick666

OpenStudy (fibonaccichick666):

yes, so what does doing that give you?

OpenStudy (anonymous):

oops i was computing x=3*(2y+3) the answer to yours is actually y=(1/2)(x^3-3)

OpenStudy (moonlight93):

x=(2y+3)^3 @FibonacciChick666 do I multiply the (2y+3) or is there an easier way to solve for y

OpenStudy (fibonaccichick666):

well, so let's talk about "x=(2y+3)^3"

OpenStudy (anonymous):

x=(2y+3)^(1/3) x^3=2y+3 2y=x^3-3 y=(1/2)(x^3-3)

OpenStudy (fibonaccichick666):

when you set both sides to the third power, you actually will get this: \[(x)^3=[(2y+3)^\frac{1}{3}]^3\]

OpenStudy (fibonaccichick666):

can you simplify that for me?

OpenStudy (moonlight93):

8 y^3+36 y^2+54 y+27

OpenStudy (fibonaccichick666):

how did you get that?

OpenStudy (moonlight93):

(2y+3)^3

OpenStudy (fibonaccichick666):

ok, so take a step back, if you have \[(x^m)^n\] how do you simplify that?

OpenStudy (moonlight93):

x^mn

OpenStudy (fibonaccichick666):

ok you multiply them right?

OpenStudy (moonlight93):

yes

OpenStudy (fibonaccichick666):

so go back and take a look at what I wrote

OpenStudy (fibonaccichick666):

here:\[ (x)^3=[(2y+3)^\frac{1}{3}]^3\]

OpenStudy (fibonaccichick666):

so take a closer look at that right side

OpenStudy (moonlight93):

x = (2 y+3)^(1/3)

OpenStudy (fibonaccichick666):

no, remember, how do you simplify \[(x^m)^n\]

OpenStudy (fibonaccichick666):

can you identify your m and n?

OpenStudy (moonlight93):

x^3= 2y+3

OpenStudy (moonlight93):

1/3 times 3 is 1

OpenStudy (fibonaccichick666):

ahh there we go, ok so you have \[x^3=2y+3\]

OpenStudy (fibonaccichick666):

so, how can you get y by itself?

OpenStudy (fibonaccichick666):

step by step please

OpenStudy (moonlight93):

switch it to the other side and change its sign -2y= 3-x^3

OpenStudy (moonlight93):

y = (x^3/2 ) - (3/2)

OpenStudy (moonlight93):

@FibonacciChick666 is this the answer?

OpenStudy (fibonaccichick666):

you have to remember to divide by negative 2 if you are going to take the step to put y on the left

OpenStudy (fibonaccichick666):

(you do not have to do that step, y can stay on the right side)

OpenStudy (fibonaccichick666):

ah wait you did, i misread, my bad

OpenStudy (moonlight93):

x^3-3=2y

OpenStudy (fibonaccichick666):

yes, either way yields the same answer, you were correct. I was thinking of this way when I saw the placement

OpenStudy (moonlight93):

thank you so much for helping me, :) @FibonacciChick666

OpenStudy (fibonaccichick666):

np, thanks for upholding the code of conduct :)

OpenStudy (moonlight93):

^_^